OpenStudy (anonymous):

If f(x) = x² + 1 and g(x) = 3x + 1, find f(-1) + 5g(-3).

OpenStudy (campbell_st):

well f(-1) is \[f(-1) = (-1)^2 + 1\] and 5g(-3) is \[5g(-3) = 5 \times (3 \times(-3) + 1)\] so f(-1) + 5g(-3) is \[(-1)^2 + 1 + 5\times(3 \times (-3) + 1)\] just evaluate it

OpenStudy (anonymous):

-38?

OpenStudy (campbell_st):

seems correct to me